# Causal Discovery in Financial Markets: A Framework for Nonstationary Time-Series Data

### v2

@@ Abstract
A deeper comprehension of financial markets necessitates understanding not only the statistical dependencies among various entities but also the causal dependencies. This paper extends the Constraint-based Causal Discovery from Heterogeneous Data algorithm to account for lagged relationships in time-series data (an algorithm we call CD-NOTS), shedding light on the complex causal relations between different financial assets and variables. We compare the performance of different algorithmic choices, such as the choice of conditional independence test, to give general advice on the effective way to use CD-NOTS. Using the results from the simulated data, we apply CD-NOTS to a broad range of indices and factors in order to identify causal connections among the entities, thereby showing how causal discovery can serve as a valuable tool for factor-based investing, portfolio diversification, and comprehension of market dynamics. Further, we show our algorithm is a more effective alternative to other causal discovery algorithms since the assumptions of our algorithm are more realistic in terms of financial data, a conclusion we find is statistically significant.
@@ 1 Introduction
Actions have consequences; that explains causality in simple words. Going beyond correlation or association, it is a guide which helps us gain a deeper understanding of the relationships between events, allowing us to perceive the impact of one phenomenon on another. An empirical causal framework would consist of three phases: (a) causal discovery; (b) causal inference; (c) causal explainability. In the first phase, the causal network is estimated from the data based on independence tests. When the network is discovered, then different scenarios can be tested on the structure. Afterwards, we can dive deep into the intuitive why of the causal network and results.
In this paper, our focus is on causal discovery, which aims to uncover the causal dependency structure from the observed data (Spirtes et al., 2001). While the traits in the data narrow down the possible causal network structures, it may not necessarily be sufficient to discover the complete underlying true causal network.
@@
Most causal discovery methods can be categorized into three types of algorithms: constraint-based (e.g., Spirtes et al. (2000); Huang et al. (2020)), score-based (e.g., Chickering (2003); Silander and Myllymäki (2012); Lam et al. (2022)), and functional causal model-based (FCM) (e.g., Shimizu et al. (2006, 2011); Hyvärinen et al. (2010)). The constraint-based causal discovery algorithm infers causal dependency between nodes by running statistical independence tests among conditional and marginal probabilities to establish the network structure. The score-based algorithm defines a score criterion (e.g., likelihood scores Schwarz (1978); Buntine (1991)) and optimizes for the causal structure that best fits the data. On the other hand, FCMs intend to model the functional mechanisms of the variables. This calls for additional assumptions on the distribution of the data or domain-specific knowledge.
@@
Causal discovery algorithms require assumptions about the data generating process; thus, each algorithm has limitations depending on the realism of the assumptions. For example, some only work on linear systems (Granger, 1969), some assume a Gaussian distribution for the variables (Ramsey and Andrews, 2017), and many assume there are no latent (unobserved) confounders (Zhang and Hyvärinen, 2009). Further, implementing causal discovery algorithms on time-series data adds additional challenges, such as sampling interval and nonstationarity, to the existing issues. There are ways to reduce the nonstationarity of time-series data such as using stock returns as opposed to price. However, empirically, we find this is insufficient to make the process stationary. Unlike evaluation of predictive performance, in general, causal discovery algorithms, when applied in practice, do not have ground truth data or ways to test the accuracy of the discovered network via randomized controlled trials. This leads to the importance of ensuring that the assumptions of the causal discovery algorithm pertains to the situations in which they are used. In our case studies (Section 5), we show that, in the financial examples we experiment with, the data is neither linear, Gaussian, nor stationary.
Refer to captionptCD-NOTSRefer to caption
@@
Figure 1: We show a high-level diagram of the problem we are solving. We insert some time-series data into our causal discovery algorithm, from which we get a causal diagram of how the variables are related.
Our developed algorithm, Constraint-based Causal Discovery from Heterogeneous Time-Series (CD-NOTS), overcomes the limitations mentioned above. It is a nonparametric approach, meaning it can capture both linear and non-linear causal relations and can also capture non-Gaussian distributions. It is capable of handling nonstationarity, and it is able to detect both lagged and contemporaneous causalities. Prior work has developed and used causal discovery algorithms that capture subsets of these properties; our work is the first (to the best of our knowledge) to tackle all three. We then analyze the patterns found in financial data using this algorithm.
@@
The structure of this paper is as follows. In Section 2, we outline the related work and literature. Moving on to Section 3, we explain our methodology in detail. In Section 4, we conduct an investigation of simulated and real financial data to assess the framework’s validity and performance. Finally, in Section 5, we summarize the key findings of the paper and propose potential avenues for future research.
The most similar to our work is Huang et al. (2020), which we extended to handle lagged relationships in time-series data. In Section 3.3, we show a simple example in which we illustrate the difference in the causal relationships discovered when accounting for lagged relationships. Huang et al. (2020) discusses in their work a way in which their algorithm can be extended to lagged relationships; however, in Section 3.4, we show a simple example in which their extension would discover an incorrect graph.
2 Related Work
Causality explores the mechanisms that influence the outcomes, providing a more comprehensive picture of the why and how of unfolding events. The Rubin Causal Model (Rubin, 1974) and Pearl’s Structural Causal Model (SCM) (Pearl, 2009) are two approaches that are mainly considered when talking about causal relations.. A Structural Causal Model (Pearl, 2014) defines the causal relationships between variables via functions, taking the cause and the effect variables as input and output respectively. The SCM can be represented as a graphical model which can then be used for causal inference. On the other hand, the Rubin model (Splawa-Neyman et al., 1990), also known as the Potential Outcomes Framework, defines the conditional independence properties of the treatment assignment mechanism in order to analyze the unit-based response variable. Often, this framework requires a set of covariates that satisfy conditional ignorability, or in other words, no unobserved confounders. Note that prior work has shown that SCMs are formally equivalent to the Potential Outcomes Framework (Pearl, 2009).
Many statisticians have formulated multiple computational techniques and procedures to discover causal links through learning the causal relationships, typically depicted as directed acyclic graphs, as reviewed in Glymour et al. (2019). After obtaining the causal skeleton, there is an attempt to orient the edges. The two main methods are constraint-based and score-based. Constraint-based ones can handle confounders as the score-based approach is applicable only when there is a no-confounder assumption.
Constraint-based algorithms run independence-based tests to establish the underlying causal network (Nogueira et al., 2022). This set of algorithms starts from a full graph, orients edges based on conditional independence (CI), and identifies colliders. It is an iterative process which ends when there are no further edges to orient. Two widely used algorithms are Peters and Clark (PC), and Fast Causal Inference (FCI) (Spirtes et al., 2000). PC assumes the absence of confounders, which results in asymptotically correct causal information that helps with simplifying the problem. There have been improvements to the PC algorithm that tend to reduce the number of conditional independence tests and relax the assumptions (e.g., Tsamardinos et al. (2006); Bühlmann et al. (2009); Ramsey et al. (2012)). On the other hand, FCI produces asymptotically correct outcomes even when confounders are present; Colombo et al. (2012) and Rohekar et al. (2021) are two improved FCI works. Since our work is an extension of the PC algorithm, we utilize four different tests: Kernel-based Conditional Independence Test (KCIT) (Zhang et al., 2012), Randomized Conditional Correlation Test (RCoT) (Strobl et al., 2017), Conditional Mutual Information K-nearest neighbor (CMIknn) Test (Runge, 2017), and Partial Correlation (ParCorr) (Baba et al., 2004).
More recently, causal structure discovery algorithms have been developed that do not require statistical tests but instead use assumptions of the functional form or of the distribution in order to recover the structure (Shimizu et al., 2006; Zhang and Hyvärinen, 2009, 2010). For example, these algorithms assume absence of confounders and a non-Gaussian distribution within the data, enabling them to uncover the asymmetry inherent in the causes and effects distribution (Gendron et al., 2023).
Causal discovery in time-series data is about uncovering and quantifying causal relationships among variables that evolve over time. The difference between causal discovery methods and traditional time-series analysis is that causal discovery aims to identify the cause and effect relationships in the data, while traditional time-series analysis is more focused on prediction.
Granger causality (Granger, 1969) is an approach that estimates the linear causal relations using a VAR model (Hyvärinen et al., 2010). There is additional work for causal structure recovery using Dynamic Bayesian Networks and Hidden Markov models (Moraffah et al., 2021). Another class of causal discovery methods uses information theoretic quantities such as Mutual Information (Palus et al., 2001), and Transfer Entropy (Schreiber, 2000). Furthermore, graphical approaches like Time-Series Fast Causal Inference (Entner and Hoyer, 2010) and PCMCI (Runge et al., 2019) are more recent graphical approaches that are tweaked to be able to handle time-series data.
Most of the time-series causal discovery literature mentioned are focused on stationary time-series, and their approaches might lead to spurious causal relations if the time-series contains distributional shifts. Some work has been developed to address this gap (e.g., Peters et al. (2015); Zhang et al. (2017)). One approach is to assume the consistency of the nonstationarity reflected in the network across time. For instance, Huang et al. (2020) proposes an algorithm for heterogeneous time-series that only checks the nonstationarity at one time $t$. Our developed algorithm extends this to handle nonstationarity in lagged relationships.
@@ 3 Methodology
Time-series is the fundamental type of data used in many domains, from finance to climate science and engineering, requiring customizing algorithms to take this type of data as input. This tailoring requires taking into account some of the characteristics that this type of data carries, such as temporality, lead-lag dependency, and nonstationarity. Similar to CD-NOD, our algorithm, Constraint-based Causal Discovery from Heterogeneous Time-Series (CD-NOTS), has four steps: adding the time-indexed node, using conditional independence tests to discover the causal skeleton (undirected causal graph), orienting the edges via prior knowledge and specific structures (e.g., the arrow of time and V-structures), and, finally, orienting remaining edges based on causal change independence. By design, our framework is able to handle time-series as input and, accordingly, takes into account its unique characteristics; in comparison, CD-NOD does not handle lagged relationships (Section 3.4).
With suitable assumptions, the causal network, using only observational data, can be estimated using conditional independence tests (Schölkopf et al., 2021). We follow the same set of assumptions commonly used in the causal network field.
Assumption 1 (Causal Sufficiency).
We assume that any potential confounders can be expressed as a smooth function of time. Hence, the time index node is a representation of all confounders.
Assumption 2 (Causal Faithfulness).
@@
Assumption 4 (Causal Consistency).
We assume that the causal relations between the variables, including the time indexed node, are consistent through time. This means that when lagged relations are present, the contemporaneous causal relations are the same in each lag and the cross lag causal relations also repeat themselves.
@@ 3.1 CD-NOTS Algorithm
We extend the algorithm described in Huang et al. (2020) in order to identify the causal network for time-series. Our developed approach, termed Constraint-based Causal Discovery from Nonstationary Time-Series (CD-NOTS), enhances the original method for identifying causal relationships between time-series. We show a high-level diagram of our algorithm in Figure 2.
@@
Assume we have $N$ time-series. Within our network setup, each time-series is treated as a distinct node, denoted as $V_{i,t}$, representing the time-series $i$ at time $t$. To handle nonstationarity, we introduce a time-indexed node, $U_{t}$. Specifically, for any nonstationary time-series $i$, we anticipate a connection between $V_{i,t}$ and $U_{t}$ across all time points. We start from a full graph including the time-indexed node, where all the edges are drawn out.
@@
With Assumptions 1, 2 and 3 in place, we reconstruct the causal skeleton. Employing CI tests, we investigate the independence between a node and the time-indexed node given a subset of nodes $S$. This process figures out the edges between nodes $U_{t}$ and $V_{i,t}$. The hypothesis test is:
$H_{0}\mathrel{\mathop{:}}P(V_{i,t},U_{t}|S)=P(V_{i,t}|S)\,P(U_{t}|S)\,,$
$H_{1}\mathrel{\mathop{:}}P(V_{i,t},U_{t}|S)\neq P(V_{i,t}|S)\,P(U_{t}|S)\,,$
@@
Afterwards, utilizing the same test, we evaluate the independence of node pairs $V_{i,t-l}$ and $V_{j,t}$ excluding $U_{t}$. The null hypothesis is that the pair are conditionally independent, implying that their conditional probability equals the product of their marginals. If the obtained p-value falls below the confidence level, signifying dependence, the edge between the pair remains. The hypothesis test is:
$H_{0}\mathrel{\mathop{:}}P(V_{i,t-l},V_{j,t}|S)=P(V_{i,t-l}|S)\,P(V_{j,t}|S)\,,$
$H_{1}\mathrel{\mathop{:}}P(V_{i,t-l},V_{j,t}|S)\neq P(V_{i,t-l}|S)\,P(V_{j,t}|S)\,,$
where $S\subseteq\{U_{t}\}\cup\{V_{k,t-m}\ |\ k=1,...,N;m=0,...,L\}\setminus\{V_{i,t-l},V_{j,t}\}$. It is worth noting that when $l=0$, indicating no lag, we capture the contemporaneous relationships between nodes. On the other hand, when $l>0$, we discover the lagged dependencies among the time-series. Under Assumption 4, we maintain the structure consistent over time. This means that the discovered dependencies between $V_{i,t-l},V_{j,t}$ remain the same for $V_{i,t-l-m},V_{j,t-m};m=1,...,l-1$. The discovered network is undirected and has the edge minimality condition. This condition signifies that the network has the smallest number of causal changing mechanisms. This is due to the faithfulness across the entire network (Ghassami et al., 2018).
@@
Once we have derived the network skeleton, our next step involves orienting the edges based on existing knowledge. If there are any nonstationary nodes, meaning there exists an edge between the node and time indexed node, then the orientation is from the time indexed node to the nonstationary node, as time is what causes the node’s corresponding time-series to be nonstationary. Moreover, all the edges where $l<m$ between pairs $V_{i,t-l}$ and $V_{j,t-m}$, the orientation is from $V_{i,t-m}$ to $V_{j,t-l}$, as logically the effect would happen after the cause. These directed edges along with the conditional independence tests can be used to orient other edges using the unique properties of V-structures (colliders) (Spirtes et al., 2000; Meek, 1995).
@@
There might still remain some edges that are not yet directed. In the fourth stage, we rely on the independent changes of the causal modules (Pearl, 2009). A dependency measure is defined as an extension of the Hilbert Schmidt Independence Criterion (Gretton et al., 2008), which shows the level of dependency between causal modules. This is done by developing a kernel embedding of nonstationary conditional distributions and using their Gram matrix to create a test statistic (Huang et al., 2017). We apply the Meek orientation rule (Meek, 1995) on top of this to orient other edges. Note that there may remain edges that are undirected even after the third and fourth stage; in this scenario, no comments can be given on the causal direction of the corresponding pair of nodes.
@@
3.2 Testing the Assumptions
The main assumptions we claim that are incorrectly assumed in prior causal discovery algorithms is stationarity, no lagged relationships, and linearity. Given the fact our algorithm does not make the first two assumptions, we can analyze the graph from our algorithm to see if nonstationary and lagged relationships are found; if found, then the conclusion is that, assuming stationarity (lagged relationships) is the null hypothesis, the statistical tests could not reject the null hypothesis.

### v3

@@ Abstract
This paper introduces a new causal structure learning method for nonstationary time series data, a common data type found in fields such as finance, economics, healthcare, and environmental science. Our work builds upon the constraint-based causal discovery from nonstationary data algorithm (CD-NOD). We introduce a refined version (CD-NOTS) which is designed specifically to account for lagged dependencies in time series data. We compare the performance of different algorithmic choices, such as the type of conditional independence test and the significance level, to help select the best hyperparameters given various scenarios of sample size, problem dimensionality, and availability of computational resources. Using the results from the simulated data, we apply CD-NOTS to a broad range of real-world financial applications in order to identify causal connections among nonstationary time series data, thereby illustrating applications in factor-based investing, portfolio diversification, and comprehension of market dynamics.
Keywords: Causal discovery, Structure learning, Nonstationary, Time series, Factor investing
@@ 1 Introduction
Actions have consequences; that explains causality in simple words. Going beyond correlation or association, it is a guide which helps us gain a deeper understanding of the relationships between events, allowing to perceive the impact of one phenomenon on another. An empirical causal framework would consist of three phases: (a) causal discovery; (b) causal inference; (c) causal explainability. In the first phase, the causal network is estimated from the data based on independence tests. When the network is discovered, then different scenarios can be tested on the structure. Afterwards, we can dive deep into the why and explanations of the scenarios implied by the discovered network.
In this paper, our focus is on causal discovery in time series, which aims to uncover the causal dependency structure from the observed data (Figure 1). Time series is the fundamental type of data used in many domains, from finance to climate science and engineering, requiring customized algorithms to take into account characteristics such as temporality, lead-lag dependency, and nonstationarity.
@@
Most causal discovery methods can be categorized into three types of algorithms: constraint-based (e.g., Spirtes et al. (2000); Huang et al. (2020)), score-based (e.g., Chickering (2003); Silander and Myllymäki (2012); Lam et al. (2022)), and functional causal model-based (FCM) (e.g., Shimizu et al. (2006, 2011); Hyvärinen et al. (2010)). The constraint-based causal discovery algorithm infers causal dependency between nodes by running statistical independence tests among conditional and marginal probabilities to establish the network structure (Tsagris et al., 2018; Nogueira et al., 2022). Two widely used algorithms of this type are Peters and Clark (PC), and Fast Causal Inference (FCI) (Spirtes et al., 2000). The score-based algorithm defines a score criterion (e.g., likelihood scores (Schwarz, 1978; Buntine, 1991)) and optimizes for the causal structure that best fits the data. On the other hand, FCMs intend to model the functional mechanisms of the variables. This calls for additional assumptions on the distribution of the data or domain-specific knowledge.
@@
Causal discovery algorithms require assumptions about the data generating process; thus, each algorithm has limitations depending on the realism of the assumptions. For example, some only work on linear systems (Granger, 1969), some assume a Gaussian distribution for the variables (Ramsey and Andrews, 2017), and many assume there are no latent (unobserved) confounders [^1] Some research has been conducted to first eliminate the impact of hidden variables and then apply the structural learning model (e.g., Frot et al. (2019))..(Zhang and Hyvärinen, 2009). Further, implementing causal discovery algorithms on time series data adds additional challenges, such as nonstationarity, to the existing issues. There are ways to reduce the nonstationarity of time series data such as using stock returns as opposed to prices. However, empirically, we find this is insufficient to make the process stationary. Unlike evaluation of predictive performance, in general, causal discovery algorithms, when applied in practice, do not have ground truth data or ways to test the accuracy of the discovered network via randomized controlled trials. This leads to the importance of ensuring that the assumptions of the causal discovery algorithm pertain to the situations in which they are used. In our case studies (Section 4), we show that, in the financial examples we experiment with, the data is neither linear, Gaussian, nor stationary.
There is a line of approaches addressing the time series causal discovery problem such as Granger causality (Granger, 1969), Dynamic Bayesian Networks and Hidden Markov models (Moraffah et al., 2021), Mutual Information (Palus et al., 2001), Transfer Entropy (Schreiber, 2000), time series Fast Causal Inference (Entner and Hoyer, 2010), and PCMCI (Runge et al., 2019). Since most of the literature mentioned are focused on stationary time series and most current approaches do not handle heterogeneous variable types their approaches might lead to spurious causal relations if the time series contains distributional shifts. Some work has been developed to address this gap (e.g., Peters et al. (2015); Zhang et al. (2017)).
Refer to captionCD-NOTSRefer to caption
@@
Figure 1: High-level diagram of the problem; we insert some time series data into our causal discovery algorithm, from which we get a causal diagram of how the variables are related.
We design an algorithm for causal discovery from nonstationary time series, referred to as the “Constraint-based Causal Discovery from Nonstationary Time Series (CD-NOTS)” hereafter. This algorithm is an extension of the CD-NOD algorithm (Huang et al., 2020) to handle time series data [^2] Huang et al. (2020) discusses in their work a way in which their algorithm can be extended to lagged relationships; however, in Section 2.4, we show a simple example in which their extension would discover an incorrect graph.. CD-NOTS handles nonstationarity as well as pseudo-confounders (defined in Section 2) while considering lagged relationships and enforces consistency of the graph through time. It is a nonparametric approach, meaning it can capture both linear and non-linear causal relations and also work with non-Gaussian distributions (e.g., Wang and Drton (2023)). In addition, it is able to detect both lagged and contemporaneous causal effects. The algorithm starts from a full graph, considering nodes for lagged time series, and designates a node to the time itself in order to model nonstationarity. Then, it uses Conditional Independence (CI) tests to uncover the causal skeleton. Leveraging background knowledge (e.g., Andrews et al. (2023)), discoverable structures, and causal change independence, it orients the edges. We experiment CD-NOTS with four different tests: Kernel-based Conditional Independence Test (KCIT) (Zhang et al., 2012), Randomized Conditional Correlation Test (RCoT) (Strobl et al., 2017), Conditional Mutual Information K-nearest neighbor (CMIknn) Test (Runge, 2017), and Partial Correlation (ParCorr) (Baba et al., 2004).
@@
The structure of this paper is as follows. In Section 2, we explain our methodology in detail. In Section 3, we conduct an investigation of simulated and real financial data to assess the framework’s validity and performance. Finally, in Section 4, we summarize the key findings of the paper and propose potential avenues for future research.
@@ 2 Methodology
Following the CD-NOD approach, our algorithm, CD-NOTS, consists of four steps: adding the time-indexed node (which helps in handling nonstationarity), discovering the causal skeleton (undirected causal graph) using conditional independence tests, orienting the edges via prior knowledge (e.g., the arrow of time) and identifiable structures (e.g., V-structures) and, finally, orienting remaining edges based on causal change independence.
With suitable assumptions, the causal network using only observational data, can be estimated by conditional independence tests (Schölkopf et al., 2021). We follow the same set of assumptions commonly used in the causal network field.
Assumption 1 (Pseudo Causal Sufficiency)
We assume that any potential confounders can be expressed as a smooth function of time. Hence, the time-indexed node is the representative of all the confounders.
Assumption 2 (Causal Faithfulness)
@@
Assumption 4 (Causal Consistency)
We assume that the causal relations between the variables, including the time-indexed node, are consistent through time. This means that when lagged relations are present, the contemporaneous causal relations are the same in each lag and the cross lag causal relations repeat themselves.
We note that most prior work assumes sufficiency, i.e., cannot handle any unobserved confounders.
@@ 2.1 CD-NOTS Algorithm
Our developed approach, CD-NOTS, identifies causal relationships between time series. We show a high-level diagram of our algorithm in Figure 2.
@@
Assume we have $N$ time series. Within our network setup, each time series is treated as a distinct node, denoted as $V_{i,t}$, representing the time series $i$ at time $t$. To handle nonstationarity, we introduce a time-indexed node, $U_{t}$. Specifically, for any nonstationary time series $i$, we anticipate a connection between $V_{i,t}$ and $U_{t}$ across all time points. We start from a full graph including the time-indexed node, where all the edges are drawn out.
@@
With Assumptions 1, 2 and 3 in place, we reconstruct the causal skeleton. Employing CI tests, we investigate the independence between a node and the time-indexed node given a subset of nodes $S$. This process figures out the edges between nodes $U_{t}$ and $V_{i,t}$. The null hypothesis is that the pair are conditionally independent, implying that their conditional joint probability equals the product of their conditional marginals. If the obtained p-value falls below the user specified confidence level $\alpha$, signifying dependence, the edge between the pair remains. The hypothesis test is:
$H_{0}:P(V_{i,t},U_{t}|S)=P(V_{i,t}|S)\,P(U_{t}|S)\,,$
$H_{1}:P(V_{i,t},U_{t}|S)\neq P(V_{i,t}|S)\,P(U_{t}|S)\,,$
@@
Afterwards, utilizing the same test, we evaluate the independence of node pairs $V_{i,t-l}$ and $V_{j,t}$ excluding $U_{t}$. The hypothesis test is:
$H_{0}:P(V_{i,t-l},V_{j,t}|S)=P(V_{i,t-l}|S)\,P(V_{j,t}|S)\,,$
$H_{1}:P(V_{i,t-l},V_{j,t}|S)\neq P(V_{i,t-l}|S)\,P(V_{j,t}|S)\,,$
where $S\subseteq\{U_{t}\}\cup\{V_{k,t-m}\ |\ k=1,...,N;m=0,...,L\}\setminus\{V_{i,t-l},V_{j,t}\}$. It is worth noting that when $l=0$, indicating no lag, we capture the contemporaneous relationships between nodes. On the other hand, when $l>0$, we discover the lagged dependencies among the time series. Under Assumption 4, we maintain the structure consistent over time. This means that the discovered dependencies between $V_{i,t-l},V_{j,t}$ remain the same for $V_{i,t-l-m},V_{j,t-m};m=1,...,L-l$. The discovered network is undirected and has the edge minimality condition. This condition signifies that the network has the smallest number of causal changing mechanisms. This is due to the faithfulness across the entire network (Ghassami et al., 2018).
@@
Once we have derived the network skeleton, our next step involves orienting the edges based on prior knowledge and identifiable structures. If there are any nonstationary nodes, meaning there exists an edge between the node and the time-indexed node, then the orientation is from the time-indexed node to the nonstationary node, as time is what causes the node’s corresponding time series to be nonstationary. Moreover, all the edges where $l<m$ between pairs $V_{i,t-l}$ and $V_{j,t-m}$, the orientation is from $V_{i,t-m}$ to $V_{j,t-l}$, as logically the effect would happen after the cause. These directed edges along with the conditional independence tests can be used to orient other edges using the unique properties of V-structures (colliders) (Spirtes et al., 2000; Meek, 1995).
@@
There might still remain some edges that are not yet directed. In the fourth stage, we rely on the independent changes of the causal modules (Pearl, 2009). A dependency measure is defined as an extension of the Hilbert Schmidt Independence Criterion (Gretton et al., 2007), which shows the level of dependency between causal modules. This is done by developing a kernel embedding of nonstationary conditional distributions and using their Gram matrix to create a test statistic (Huang et al., 2017). We apply the Meek orientation rule (Meek, 1995) on top of this to orient other edges. Note that this stage is only applicable to the edges between two nonstationary nodes. There may remain edges that are undirected even after the third and fourth stages; in this scenario, no comments can be given on the causal direction of the corresponding pair of nodes.
@@
2.2 Testing Assumptions
In practice, an important step when using algorithms that learn from data is to evaluate (or validate) the results of the algorithm. Running interventional studies is one approach; however, this is often infeasible, if not impossible, in fields such as finance. Instead, as mentioned in the introduction, evaluation of causal discovery algorithms in the practice rests upon evaluating the realism of the assumptions underlying the algorithm.
We categorize assumptions into three types:
  1. 1.
    Objective Assumptions: These are assumptions which are statistically testable such as stationarity and linearity. In Section 2.2.1, we discuss specific ways we can test these two properties. Since these assumptions are testable, we argue that any causal discovery algorithms reliant on these assumptions should ensure that the data aligns with them.
  2. 2.
    Algorithmic Assumptions (or Algorithmic Hyperparameters): These assumptions refer to decisions in the causal discovery algorithm which can neither be directly evaluated nor justified purely by domain expertise. Examples include the significance threshold used in the conditional independence test and the type of conditional independence test. While there is some theories and studies that can be used to justify one threshold or test over another (e.g., our simulated experiments in Section 3), we recommend testing the robustness of the causal discovery algorithm’s results to these hyperparameters.
  3. 3.
    Subjective Assumptions: These are assumptions which cannot be statistically tested and require domain expertise to evaluate. Pseudo-sufficiency, consistency, faithfulness, and randomness (assumptions made by CD-NOTS) fall in this category. In our case studies, while we maintain confidence in the validity of these assumptions, the credibility of our algorithm’s results comes from the reader’s belief in these assumptions. We discuss the intuition behind these assumptions in Section 2.2.2.
2.2.1 Objective Assumptions
One of the main assumptions in prior causal discovery algorithms is stationarity, which can result in serious inaccuracies when dealing with inherently nonstationary data, such as those commonly encountered in fiance and economics. Since our algorithm does not rely on this assumption, we can analyze the graph identified by our algorithm to see if nonstationary relationships are found.